Let’s say that, in June of 2006, a national park sees 20 fires begin. If fire-starts are random, what is the probability that: a. on a given day, NO fires will start. b. on a given day, two or more fires will start.
Answer #1

since June has 30 days , expected number of fires in a day =20/30 =2/3

a)

from Poisson approximation:

P(on a given day, NO fires will start) =P(X=0)=e-2/3*(2/3)0/0! =0.5134

b)

P(on a given day, two or more fires will start) =P(X>=2)=1-P(X<=1)

=1-(P(X=0)+P(X=1))=1-(e-2/3*(2/3)0/0!+e-2/3*(2/3)1/1!) =0.1443